Shift report 20011031 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Oct 31 2001 - 23:52:13 EST

  • Next message: Apache: "Shift report 20011101 00:00-08:00"

    ------------------------------------------------------------
    Date: Wed Oct 31 23:52:12 2001, Shift: 16:00-24:00
    Supervisor: RW, with: 
    Log book#10 pages 79 - 80
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam came back at 17:30 and physics run started at 17:55.
    Blue/Yellow=19.5/22.0 ZDC rate 350. At ZDC rate of 120, DUMP
    was requested by BRAHMS and PHENIX. Beam was dumped at 20:55.
    At 22:30 a very good storage was made. Blue/yellow 28.0/28.0 
    ZDC 1200. (STAR & PHENIX about 2000) No increase by an 
    additional steering request. 
    
    2. Run 5601 TPC Pedestal run
           5602 DVM run
           5603 Pedestals
           5604 Data FS@12deg. 1/4A  MRS @45deg 4A
    	    Trigger 6 105K
           5605 Trigger 6 44K, stopped by dumping beam.
           5606 Trigger 6 100K start with new storage.
           5607 still running.
    
    3. T5 Mod 2 still off. No other problems are detected.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Oct 31 2001 - 23:52:36 EST